OpenStudy (anonymous):

how is tan (3pi/4) = -1?

OpenStudy (anonymous):

You can think of it geometrically. First if you notice tan(3pi/4) = -tan(pi/4) you can draw a right angled triangle with one of its angles equal to pi/4. Then look at the ratio of the side opposite to that angle to the one adjacent as this is why tan really is. Draw it out and you should understand your result geometrically.

OpenStudy (anonymous):

could you explain this in mathematical way? as in like what equals to sin x and what equals to cos x and when you divide those two you get -1?

OpenStudy (anonymous):

Geomtry is mathematical :) tanx = sinx / cosx

OpenStudy (anonymous):

as in like numbers?

OpenStudy (anonymous):

isn't sin = root 3 /2 and cos = 1/2?

OpenStudy (anonymous):

and when you divide, you get root 3

OpenStudy (anonymous):

i think i'm doing wrong there..

OpenStudy (anonymous):

oh wait i get it now ty

OpenStudy (anonymous):

sinx = sqrt(2) / 2 cosx = -sqrt(2) / 2 sinx / cosx = -1

OpenStudy (anonymous):

got it, thanks!